What is the mass of
  0.2 mole of oxygen atoms?
What is the mass of
  0.2 mole of oxygen atoms?

Grade:12th pass

2 Answers

Harshit Singh
3 years ago
Dear Student

Mole of Oxygen atoms = 0.2 mole
Molar mass of oxygen atoms = 16 g
Mass of oxygen atoms = 16 × 0.2 = 3.2 g

Thanks
Shraddha Singh
3 years ago
Dear Student,
 
Molar Mass of Oxygen = 16 g
Moles of Oxygen = 0.2
Mass of Oxygen = 16 × 0.2 = 3.2 g
 
Hope it helps!
Thank you ✌️